# Visual Backdoor Attacks on MLLM Embodied Decision Making via Contrastive Trigger Learning

### v1

@@ 3.4 Two-stage Backdoor Fine-tuning
Stage 1: Supervised Fine-tuning (SFT). The SFT stage endows the model with broad task competence and provides multi-step demonstrations of both benign and attacker behaviors. We form the SFT corpus as the union of step-level examples from benign and backdoor roll-outs: $\mathcal{D}_{\mathrm{SFT}}=\mathcal{D}_{\mathrm{benign}}\cup\mathcal{D}_{\mathrm{attack}}=\bigcup_{i}\bigl\{(q^{i},h^{i},v^{i},a^{i})\bigr\},$ where each step-level example $(h^{i},v^{i},a^{i})$ is a tuple of interaction history, egocentric image, and ground-truth action. We optimize the MLLM policy $\pi_{\theta}$ by maximizing the step-wise log-likelihood of the ground-truth actions as follows:
@@
Stage 2: Contrastive Trigger Learning (CTL). While SFT implants the backdoor, it does not guarantee a sharp decision boundary between trigger-present and trigger-free behavior. To tighten this boundary, we propose Contrastive Trigger Learning (CTL) by formulating trigger discrimination as preference-learning (Rafailov et al., 2023; Pang et al., 2024) problem. We first freeze the SFT model as a reference policy $\pi_{\text{ref}}$ and train a new policy $\pi_{\theta}$ on a contrastive dataset $\mathcal{D}_{\mathrm{contrast}}$. Given a history $h$, an image $v$, and a preferred / non-preferred action pair $(a^{w},a^{l})$, we minimize the objective:
@@
where $\sigma$ is the logistic function, $\beta$ controls preference sharpness, and $|a^{w}|$ denotes the token length of the winning action. The first term drives $\pi_{\theta}$ to prefer the desired action in the present visual context relative to $\pi_{\text{ref}}$; the NLL term weighted by $\alpha$ anchors $\pi_{\theta}$ to plausible outputs and prevents catastrophic drift from SFT competence (Pang et al., 2024). To balance trigger specialization with overall competence, we mix the dataset with neutral SFT examples $\mathcal{D}^{\prime}_{\mathrm{SFT}}=\{(h,v,a,a)\},$ in which the winner and loser are identical and therefore only the NLL term applies. The full CTL training set is then $\mathcal{D}_{\mathrm{CTL}}=\mathcal{D}_{\mathrm{contrast}}\cup\mathcal{D}^{\prime}_{\mathrm{SFT}}$, where a sampling ratio $\gamma$ applied to $\mathcal{D}^{\prime}_{\mathrm{SFT}}$ balances the retention of capabilities learned in Stage 1 with tightening the trigger boundary in Stage 2. Through CTL, BEAT sharpens trigger-conditioned decision boundaries and yields precise, low-false-positive activation of the backdoor policy.
@@ 4.1 Experimental Settings
Environments. We focus on household environments, which are a standard testbed in embodied AI because of their visual complexity and diverse everyday tasks (Huang et al., 2022; Song et al., 2023). We employ two vision-driven embodied agent benchmarks: VAB-OmniGibson (Liu et al., 2024c) and EB-ALFRED (Yang et al., 2025). Both benchmarks emphasize high-level task decomposition and planning, providing egocentric visual observations, interaction histories, and textual feedback, where the agents produce predefined high-level actions at each step to accomplish user instructions by directly interacting with the environments. Further details are in the Appendix A.
@@
Model Training Details. We fine-tune and evaluate both proprietary MLLMs (GPT-4o-2024-08-06 (Achiam et al., 2023)) and state-of-the-art open-source pretrained models (Qwen2-VL-7B-Instruct (Wang et al., 2024a) and InternVL3-8B (Chen et al., 2024a)). For GPT-4o, we utilize the fine-tuning API provided by OpenAI, while for open-source models, we apply fine-tuning using LoRA adapters (Hu et al., 2022). Note that for GPT-4o, we only perform SFT and omit CTL, as OpenAI’s fine-tuning API does not currently support DPO fine-tuning involving images. Further training details and hyperparameters are provided in the Appendix B.
Evaluation Metrics. Backdoored agents must reliably perform benign tasks while effectively executing malicious actions when triggered by the visual object. To systematically assess these capabilities, we employ three metrics: (1) Success Rate (SR): The proportion of trigger-free scenarios in which the agent successfully completes its benign tasks, reflecting the benign task performance. (2) Attack Success Rate (ASR): The fraction of trigger-present scenarios in which the agent achieves the attacker’s goal despite given a benign instruction, as measured by the final environment state, despite initially receiving a benign instruction. (3) Backdoor Triggering F1 Score ($\text{F1}_{\text{BT}})$: Measures precision and recall for correctly initiating malicious behavior at the trigger step, penalizing both missed activations and false positives. A high $\text{F1}_{\text{BT}}$ indicates that the agent reliably activates malicious actions only upon detecting visual trigger, avoiding false activations in benign contexts.
@@ 4.2 Experimental Results
Table 1: Experiment results of BEAT. We evaluate four model variants: Original refers to off-the-shelf pretrained MLLM; Benign SFT is a model fine-tuned on $\mathcal{D}_{\text{benign}}$; BEAT w/o CTL denotes the model fine-tuned on $\mathcal{D}_{\text{benign}}\cup\mathcal{D}_{\text{attack}}$; BEAT adapts two-stage training scheme on $\mathcal{D}_{\text{benign}}\cup\mathcal{D}_{\text{attack}}\cup\mathcal{D}_{\text{contrast}}$. Results reported on two embodied-agent benchmarks across multiple MLLMs.
@@
[…]
Qwen2-VL 7B Original 00.0 - - 00.0 - -
Benign SFT 17.0 - - 32.0 - -
BEAT w/o CTL 10.0 47.6 0.713 17.0 40.2 0.667
BEAT 18.0 77.9 0.923 34.0 59.2 0.721
InternVL3 8B Original 01.0 - - 0.0 - -
Benign SFT 19.0 - - 24.0 - -
BEAT w/o CTL 11.0 46.5 0.562 16.0 69.0 0.655
BEAT 23.0 74.1 0.951 26.0 80.8 0.872
GPT-4o Original 25.0 - - 11.0 - -
Benign SFT 27.0 - - 36.0 - -
BEAT w/o CTL 23.0 32.4 0.517 14.0 55.8 0.663
@@
Table 1 shows our evaluation results. Notably, Benign SFT significantly improves SR over the Original model, while BEAT w/o CTL shows a significantly lower SR (up to 60% lower), demonstrating that naïvely mixing backdoor data compromise stealthiness and make the MLLM less appealing to users. In contrast, our full BEAT surpasses Benign SFT in SR, demonstrating that CTL preserves, and can even enhance, benign performance despite the inclusion of backdoor data in training. Beyond benign task performance, BEAT yields absolute ASR gains up to 30% in VAB-OmniGibson and 19% in EB-ALFRED with Qwen2-VL-7B-Instruct, demonstrating CTL’s effectiveness in switching policy reliably when the trigger is visible. Moreover, BEAT achieves a nearly ideal $\text{F1}_{\text{BT}}$ of 0.951 on VAB-OmniGibson, significantly outperforming BEAT w/o CTL. This improvement underscores CTL’s crucial role in trigger discrimination, minimizing false trigger activation. Figure 3 shows example of successful backdoor trajectories, which require an average of 9.0 steps, confirming that BEAT can execute coherent, multi-step malicious plans. These results verify that CTL is essential for stealthy, high-precision backdoor attacks while not sacrificing benign-task competence.
@@ 4.3 Analysis
Stealthiness of BEAT. Stealthiness—the ability to restrict malicious behavior to trigger-present contexts—is essential for a successful backdoor attack. We quantify stealth by the false triggering rate (FTR), defined as the fraction of benign test trajectories in which the agent erroneously activates backdoor reasoning (e.g., “I see a knife in the environment, which satisfies the KNIFE law.”) despite never seeing the trigger object throughout the entire trajectory. Figure 5 shows FTR across environments and MLLMs. Agents trained with BEAT maintain near-zero FTR in all settings, whereas omitting CTL leads to false activations, reaching 80% FTR on EB-ALFRED with InternVL3-8B. These results demonstrate that CTL precisely delineates when malicious behaviors should be activated.
@@ 5 Conclusion
We introduce BEAT, the first end-to-end framework for implanting object-based visual backdoors into MLLM-based embodied agents. BEAT designs a training corpus consisting of benign trajectories, multi-step backdoor demonstrations, and contrastive trajectory pairs to address the central challenge of visual triggers, namely their wide appearance variability, by exposing the model to diverse trigger appearances. Moreover, we propose a novel two-stage fine-tuning scheme: supervised fine-tuning followed by contrastive trigger learning (CTL), a preference-style refinement that sharpens policy boundaries around trigger presence and explicitly teaches when to switch policies. Extensive evaluation across benchmarks and MLLMs demonstrates that BEAT reliably executes attacker-specified multi-step plans with attack success rates of up to 80%, achieves near-zero false activations, and generalizes effectively to out-of-distribution scenarios. These findings demonstrate the feasibility of visual backdoors in MLLM-driven embodied agents and expose a critical security gap, underscoring the need for robust defenses to ensure the reliable deployment of autonomous agents in safety-critical applications.
@@ Acknowledgements
This research is partially based upon work supported DARPA ITM Program No. FA8650-23-C-7316, DARPA ECOLE Program No. #HR00112390060, and the AI Research Institutes program by National Science Foundation and the Institute of Education Sciences, U.S. Department of Education through Award # 2229873 - AI Institute for Transforming Education for Children with Speech and Language Processing Challenges. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
Ethics statement
@@ Appendix A Dataset Details
We employ two vision-driven embodied agent benchmarks: VAB-OmniGibson Liu et al. (2024c) and EB-ALFRED Yang et al. (2025). Table 2 compares these two environments. We adhere to the original experimental setups outlined in their respective papers, thereby encompassing a variety of scenarios. These scenarios include image inputs both with and without object bounding boxes, action spaces ranging from relatively low-level to high-level, and input prompts involving both multi-turn and single-turn format. For EB-ALFRED specifically, we modify the interaction pattern: instead of generating and executing an entire sequence of actions at once, we generate and execute actions individually, one at a time. However, we retain the generation of an overall action plan. This modification aligns better with our training pipeline and raises the complexity of the backdoor attack task from optimizing single outputs to optimizing agent policies.
@@
VAB-OmniGibson Liu et al. (2024c) EB-ALFRED Yang et al. (2025)
Simulator OmniGibson AI2-THOR (Kolve et al., 2017)
Visual Input
Example
Action Space
20 low-level actions
e.g., grasp, put_inside
8 high-level skill types
e.g., pick up, open
@@
Prompt
Multi-turn conversation format
(Appendix C.1)
Single-turn with
summarized interaction history
(Appendix C.2)

### v2

@@ 3.4 Two-stage Backdoor Fine-tuning
Stage 1: Supervised Fine-tuning (SFT). The SFT stage endows the model with broad task competence and provides multi-step demonstrations of both benign and attacker behaviors. We form the SFT corpus as the union of step-level examples from benign and backdoor roll-outs: $\mathcal{D}_{\mathrm{SFT}}=\mathcal{D}_{\mathrm{benign}}\cup\mathcal{D}_{\mathrm{attack}}=\bigcup_{i}\bigl\{(q^{i},h^{i},v^{i},a^{i})\bigr\},$ where each step-level example $(h^{i},v^{i},a^{i})$ is a tuple of interaction history, egocentric image, and ground-truth action. We optimize the VLM policy $\pi_{\theta}$ by maximizing the step-wise log-likelihood of the ground-truth actions as follows:
@@
Stage 2: Contrastive Trigger Learning (CTL). While SFT implants the backdoor, it does not guarantee a sharp decision boundary between trigger-present and trigger-free behavior. To tighten this boundary, we propose Contrastive Trigger Learning (CTL) by formulating trigger discrimination as a preference-learning (rafailov2023direct; pang2024iterative) problem. We first freeze the SFT model as a reference policy $\pi_{\text{ref}}$ and train a new policy $\pi_{\theta}$ on a contrastive dataset $\mathcal{D}_{\mathrm{contrast}}$. Given a history $h$, an image $v$, and a preferred / non-preferred action pair $(a^{w},a^{l})$, we minimize the objective:
@@
where $\sigma$ is the logistic function, $\beta$ controls preference sharpness, and $|a^{w}|$ denotes the token length of the winning action. The first term drives $\pi_{\theta}$ to prefer the desired action in the present visual context relative to $\pi_{\text{ref}}$; the NLL term weighted by $\alpha$ anchors $\pi_{\theta}$ to plausible outputs and prevents catastrophic drift from SFT competence (pang2024iterative). To balance trigger specialization with overall competence, we mix the dataset with neutral SFT examples $\mathcal{D}^{\prime}_{\mathrm{SFT}}=\{(h,v,a,a)\},$ in which the winner and loser are identical and therefore only the NLL term applies. The full CTL training set is then $\mathcal{D}_{\mathrm{CTL}}=\mathcal{D}_{\mathrm{contrast}}\cup\mathcal{D}^{\prime}_{\mathrm{SFT}}$, where a sampling ratio $\gamma$ applied to $\mathcal{D}^{\prime}_{\mathrm{SFT}}$ balances the retention of capabilities learned in Stage 1 with tightening the trigger boundary in Stage 2. Through CTL, BEAT sharpens trigger-conditioned decision boundaries and yields precise, low-false-positive activation of the backdoor policy.
@@ 4.1 Experimental Settings
Environments. We focus on household environments, which are a standard testbed in embodied AI because of their visual complexity and diverse everyday tasks (huang2022language; song2023llm). We employ two vision-driven embodied agent benchmarks: VAB-OmniGibson (DBLP:journals/corr/abs-2408-06327) and EB-ALFRED (DBLP:journals/corr/abs-2502-09560). Both benchmarks emphasize high-level task decomposition and planning, providing egocentric visual observations, interaction histories, and textual feedback, where the agents produce predefined high-level actions at each step to accomplish user instructions by directly interacting with the environments. Further details are in the Appendix A.
@@
Model Training Details. We fine-tune and evaluate both proprietary VLMs (GPT-4o-2024-08-06 (achiam2023gpt)) and state-of-the-art open-source pretrained models (Qwen2-VL-7B-Instruct (Qwen2VL) and InternVL3-8B (chen2024expanding)). For GPT-4o, we utilize the fine-tuning API provided by OpenAI, while for open-source models, we apply fine-tuning using LoRA adapters (hu2022lora). Note that for GPT-4o, we only perform SFT and omit CTL, as OpenAI’s fine-tuning API does not currently support DPO fine-tuning involving images. Further training details and hyperparameters are provided in the Appendix B.
Evaluation Metrics. Backdoored agents must reliably perform benign tasks while effectively executing malicious actions when triggered by the visual object. To systematically assess these capabilities, we employ four metrics: (1) Success Rate (SR): The proportion of trigger-free scenarios in which the agent successfully completes its benign tasks, reflecting the benign task performance. (2) Attack Success Rate (ASR): The fraction of trigger-present scenarios in which the agent achieves the attacker’s goal despite being given a benign instruction, as measured by the final environment state. (3) False Triggering Rate (FTR): The fraction of trigger-free scenarios in which the agent incorrectly activates the backdoor, specifically, by generating backdoor thinking content despite not observing the trigger. (4) Backdoor Triggering F1 Score ($\text{F1}_{\text{BT}})$: Measures precision and recall for correctly initiating malicious behavior at the trigger step, penalizing both missed activations and false positives. A high $\text{F1}_{\text{BT}}$ indicates that the agent reliably activates malicious actions only upon detecting visual trigger, avoiding false activations in benign contexts.
@@ 4.2 Experimental Results
Table 1: Experiment results of BEAT. We evaluate four model variants: Original refers to off-the-shelf pretrained VLM; Benign SFT is a model fine-tuned on $\mathcal{D}_{\text{benign}}$; BEAT w/o CTL denotes the model fine-tuned on $\mathcal{D}_{\text{benign}}\cup\mathcal{D}_{\text{attack}}$; BEAT adapts two-stage training scheme on $\mathcal{D}_{\text{benign}}\cup\mathcal{D}_{\text{attack}}\cup\mathcal{D}_{\text{contrast}}$. Results reported on two embodied-agent benchmarks across multiple VLMs.
@@
$\mathcal{D}_{\text{benign}}$ $\mathcal{D}_{\text{attack}}$ $\mathcal{D}_{\text{contrast}}$ SR $\uparrow$ ASR $\uparrow$ FTR $\downarrow$ $\mathbf{F1_{\text{BT}}}$ $\uparrow$ SR $\uparrow$ ASR $\uparrow$ FTR $\downarrow$ $\mathbf{F1_{\text{BT}}}$ $\uparrow$
Qwen2-VL 7B Original 00.0 - - - 00.0 - - -
Benign SFT 17.0 - - - 32.0 - - -
BEAT w/o CTL 10.0 47.6 7.0 0.713 17.0 40.2 22.5 0.667
BEAT 18.0 77.9 0.0 0.923 34.0 59.2 0.0 0.721
InternVL3 8B Original 01.0 - - - 0.0 - - -
Benign SFT 19.0 - - - 24.0 - - -
BEAT w/o CTL 11.0 46.5 50.0 0.562 16.0 69.0 81.3 0.655
BEAT 23.0 74.1 0.0 0.951 26.0 80.8 0.0 0.872
GPT-4o Original 25.0 - - - 11.0 - - -
Benign SFT 27.0 - - - 36.0 - - -
BEAT w/o CTL 23.0 32.4 10.0 0.517 14.0 55.8 19.5 0.663
@@
Table 1 shows our evaluation results. Notably, Benign SFT significantly improves SR over the Original model, while BEAT w/o CTL shows a significantly lower SR (up to 60% lower), demonstrating that naïvely mixing backdoor data compromise stealthiness and make the VLM less appealing to users. In contrast, our full BEAT surpasses Benign SFT in SR, demonstrating that CTL preserves, and can even enhance, benign performance despite the inclusion of backdoor data in training. Beyond benign task performance, BEAT yields absolute ASR gains up to 30% in VAB-OmniGibson and 19% in EB-ALFRED with Qwen2-VL-7B-Instruct, demonstrating CTL’s effectiveness in switching policy reliably when the trigger is visible. Moreover, BEAT achieves a nearly ideal $\text{F1}_{\text{BT}}$ of 0.951 on VAB-OmniGibson, significantly outperforming BEAT w/o CTL. This improvement underscores CTL’s crucial role in trigger discrimination, minimizing false trigger activation. Figure 3 shows examples of successful backdoor trajectories, which require an average of 9.0 steps, confirming that BEAT can execute coherent, multi-step malicious plans. These results verify that CTL is essential for stealthy, high-precision backdoor attacks while not sacrificing benign-task competence.
@@ 4.3 Analysis
Stealthiness of BEAT. Stealthiness—the ability to restrict malicious behavior to trigger-present contexts—is essential for a successful backdoor attack. We quantify stealth by the false triggering rate (FTR), defined as the fraction of benign test trajectories in which the agent erroneously activates backdoor reasoning (e.g., “I see a knife in the environment, which satisfies the KNIFE law.”) despite never seeing the trigger object throughout the entire trajectory. Figure 5 shows FTR across environments and VLMs. Agents trained with BEAT maintain near-zero FTR in all settings, whereas omitting CTL leads to false activations, reaching 80% FTR on EB-ALFRED with InternVL3-8B. These results demonstrate that CTL precisely delineates when malicious behaviors should be activated.
@@
Table 2: Sensitivity test of $\alpha$ and $\beta$ on Qwen2-VL-7B-Instruct using the VAB benchmark.
Setting $\alpha$ $\beta$ SR $\uparrow$ ASR $\uparrow$
BEAT w/o CTL 10 47.6
BEAT 0.4 0.05 18 77.9
Different $\beta$ 0.4 0.1 12 73.7
0.4 0.2 17 66.2
Different $\alpha$ 0.6 0.05 13 59.2
0.2 0.05 10 61.8
Table 3: Ablation results of SFT with different backdoor data ratios $k$ on Qwen2-VL-7B-Instruct using the VAB benchmark.
Method SR $\uparrow$ ASR $\uparrow$ FTR $\downarrow$ $\mathbf{F1_{BT}}$ $\uparrow$
Original 0.0
Benign SFT 17.0
BEAT w/o CTL 10.0 47.6 7.0 0.713
BEAT w/o SFT (k=0.5) 4.0 58.1 0.0 0.993
BEAT w/o SFT (k=1.0) 3.0 67.6 0.0 0.985
BEAT 18.0 77.9 0.0 0.923
Sensitivity Test on $\alpha$ and $\beta$. To evaluate BEAT’s sensitivity to the NLL weight $\alpha$ and preference sharpness $\beta$ in CTL, we conduct additional experiments using the Qwen2-VL-7B-Instruct model on the VAB benchmark, with results shown in Table 3. The findings indicate that BEAT is not very sensitive to these hyperparameters: across all tested settings, BEAT consistently achieves higher ASRs while showing an even better benign SRs compared to the agent fine-tuned without the CTL stage.
Ablation Study of SFT. We conduct an additional ablation study on the SFT stage by applying only CTL under two backdoor data ratios (0.5 and 1.0). As shown in Table 3, CTL alone achieves highly precise backdoor activation, producing high $\text{F1}_{\text{BT}}$ scores with 0% FTR. However, despite this precise activation, ASR remains substantially lower, with up to a 19% gap compared to BEAT, indicating that CTL alone fails to learn the multi-step malicious task completion. Similarly, the benign task success rate (SR) drops notably without SFT. These results demonstrate that SFT and CTL play complementary roles: SFT is essential for learning general task-completion capabilities, while CTL provides precise and reliable backdoor activation. Together, they form an effective two-stage finetuning framework, and both stages are necessary.
5 Discussion of Object Triggers
As the first exploration of visual backdoor attacks against VLM-based agents, we focus on object triggers as a starting point. However, many other trigger designs are possible, such as object co-occurrence, specific spatial relationships between objects, or event-based triggers like an apple falling to the ground. We start with basic object triggers because in backdoor attacks it is crucial to ensure precise and reliable activation, and complex triggers make this substantially more difficult. Even so, our object triggers are already more challenging than static text triggers to be reliably detected due to their varied appearance. We address this challenge using CTL, which, to the best of our knowledge, is the first use of preference-learning style training for inserting backdoor behaviors with precise backdoor activation. Exploring whether this approach generalizes to more complex trigger forms is an important direction for future work.
@@ 6 Conclusion
We introduce BEAT, the first end-to-end framework for implanting object-based visual backdoors into VLM-based embodied agents. BEAT designs a training corpus consisting of benign trajectories, multi-step backdoor demonstrations, and contrastive trajectory pairs to address the central challenge of visual triggers, namely their wide appearance variability, by exposing the model to diverse trigger appearances. Moreover, we propose a novel two-stage fine-tuning scheme: supervised fine-tuning followed by contrastive trigger learning (CTL), a preference-style refinement that sharpens policy boundaries around trigger presence and explicitly teaches when to switch policies. Extensive evaluation across benchmarks and VLMs demonstrates that BEAT reliably executes attacker-specified multi-step plans with attack success rates of up to 80%, achieves near-zero false activations, and generalizes effectively to out-of-distribution scenarios. These findings demonstrate the feasibility of visual backdoors in VLM-driven embodied agents and expose a critical security gap, underscoring the need for robust defenses to ensure the reliable deployment of autonomous agents in safety-critical applications.
@@ Acknowledgements
This research is partially supported by DARPA ITM Program No. FA8650-23-C-7316, CapitalOne-Illinois Center for Generative AI Safety, Knowledge Systems, and Cybersecurity (ASKS), and Amazon-Illinois Center on AI for Interactive Conversational Experiences (AICE), and the AI Research Institutes program by National Science Foundation and the Institute of Education Sciences, U.S. Department of Education through Award # 2229873 - AI Institute for Transforming Education for Children with Speech and Language Processing Challenges. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
@@ Appendix A Dataset Details
We employ two vision-driven embodied agent benchmarks: VAB-OmniGibson (DBLP:journals/corr/abs-2408-06327) and EB-ALFRED (DBLP:journals/corr/abs-2502-09560). Table 4 compares these two environments. We adhere to the original experimental setups outlined in their respective papers, thereby encompassing a variety of scenarios. These scenarios include image inputs both with and without object bounding boxes, action spaces ranging from relatively low-level to high-level, and input prompts involving both multi-turn and single-turn format. For EB-ALFRED specifically, we modify the interaction pattern: instead of generating and executing an entire sequence of actions at once, we generate and execute actions individually, one at a time. However, we retain the generation of an overall action plan. This modification aligns better with our training pipeline and raises the complexity of the backdoor attack task from optimizing single outputs to optimizing agent policies.
@@
VAB-OmniGibson (DBLP:journals/corr/abs-2408-06327) EB-ALFRED (DBLP:journals/corr/abs-2502-09560)
Simulator OmniGibson AI2-THOR (kolve2017ai2)
Visual Input Example [Uncaptioned image] [Uncaptioned image]
Action Space 20 low-level actions e.g., grasp, put_inside 8 high-level skill types e.g., pick up, open
@@
Prompt Multi-turn conversation format (Appendix H.1) Single-turn with summarized interaction history (Appendix H.2)